quickstudent:

Please answer Right Now! A cylinder with a movable piston contains a fixed amount of gas at a constant pressure. Initially, the cylinder contains 0.25 liters of air at 0 degrees Celsius. When the temperature is increased to 35 degrees Celsius, the air will occupy what volume? 0.28 L 0.88 L 8.75 L 35.25 L

Vocaloid:

first convert both temperatures to Kelvin then use the formula V1/T1 = V2/T2 where V1 = initial volume, T1 = initial temperature, T2 = final temperature, and V2 = final volume solve for V2

quickstudent:

0.28 L